Pol is 10 years younger than Greg.In 7 years, he will be 10 years more than one half as old as Greg.How old is Greg at present ?
Answer by sachi(548)

let greg is x yrs
than pol is x-10
in 7 yrs
greg is x+7 yrs
than pol is x-10+7=x-3 yrs
pol will be 10 years more than one half as old as Greg
so x-3=10+1/2(x+7)
or 2(x-3)=20+x+7
or 2x-6=27+x
or x=27+6=33
so greg is 33 yrs
& pol is 23 yrs
ans